OpenStudy (anonymous):

what are three solutions for 5x+7y=19

OpenStudy (whpalmer4):

Pick a number for \(x\). Plug it into the equation, and find the value of \(y\) that makes the equation true. I always like to use 0 for things like this, because that usually makes the arithmetic/algebra simpler. \[5x+7y=19\]\[5(0)+7y=19\]\[7y=19\]\[y=\frac{19 } {7 }\]So one solution is \((0,\dfrac{19}{7})\) Now you know how to do it, generate 3 solutions of your own.

OpenStudy (precal):

actually there are more than 3 solutions to pick from

OpenStudy (precal):

5x+7y=19 7y=19-5x y=(19/7) - (5/7)x this is a linear function